Tohban Report 2014-10-22

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 15 Oct 2014
End Date: 22 Oct 2014
Tohban: Juan Camilo Buitrago-Casas
Tohban email:
Next Tohban: Hazel Bain
List all reports


Solar Activity

The Sun has been quiet active this week. A big active region appear on the solar disk last friday October 17. This active region hosted two X-class flares this week and a handful of M and C class flares. Currently there are five active regions on the disk. Today we have a 20% of probability to have an X-class flare and a 65% of probability for an M-class flare.

How many GOES flares occurred?

 Flares above B, C, M, X class were      0    36     9     1

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      0    26     7     1

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      0     0     0     0

There were RHESSI flares/GOES flares 410 / 46 over the time range 14-Oct-14 21-Oct-14

Memory Management

SSR was a little high this week mostly because the intensive activity of the sun and because we kept the night time taking data. We got the highest value on the SSR on 2014-10-21 around 15:00 UT, it was ~50%. Based on that we decided stopping the night time taking data today. Also, the decimation was changed to "Active/Vigorous" at 292-18:18:16 UTC.


Two warnings remained almost constantly all week regarding the Cold plate 1 [126.063 degrees (K)] and the Solar Array Temperature 1 [96.586 degrees (C)]. Some times temperature on the Cold plate 2 also appeared with a warning.

Spacecraft Operations

It looked nominal as far as a tohban can say. RHESSI decimation was changed form Normal/Vigorous to Active/Vigorous on October 19 at 18:18:16 UTC. Attenuator motion was nominal. We stopped taking the whole night time data today (October 22) at 17:40 UT. RHESSI is taking only 4 minutes on either side of the eclipse for background analysis purposes.

Data Gaps

No data gaps this week.

Detector issues

Detector 2: It still presents the light leak issue with a not significant change from last week.

There were no threshold changes this week.

Other notes

Decimation Active/Vigorous
HLAT Decimation Rear decimation weight 6, no front decimation
Night time data (fronts) Full nighttime events stopped
Night time data (rears) Full nighttime events stopped
Require extra passes? We are going to ask for extra passes this week because the big active region
Requirement for moving pointer? No
Attenuator operation Nominal with a lot of motions because the high solar activity
Detector problems? Detector 2 light leak
Personal tools